SIZING THE UBBELHODE EFFECT: THE ROTATIONAL SPECTRUM OF TERT -BUTYLALCOHOL DIMER

SHOUYUAN TANG, College of Bioengineering, ChongQing University, ChongQing, 400044, P. R. China; IRENA MAJERZ, Faculty of Chemistry, University of Wroclaw, Joliot-Curie 14, 50-383 Wroclaw, Poland; \underline{WALTHER CAMINATI}, Dipartimento di Chimica ”G. Ciamician” dell’Università, Via Selmi 2, I-40126 Bologna, Italy.

We measured the molecular beam Fourier transform microwave spectra of four isotopic species of of the dimer of tertbutanol, that is C_{4}H_{9}-OH⋯O(H)-C_{4}H_{9}, C_{4}H_{9}-OH⋯O(D)-C_{4}H_{9}, C_{4}H_{9}-OD⋯O(H)-C_{4}H_{9}, and C_{4}H_{9}-OD⋯O(D)-C_{4}H_{9}. We observed that the H – D isotopic substitution of the hydroxylic hydrogen participating in the O-H⋯O Hydrogen bond in the tert-butanol dimer produces an increase of the B and C rotational constants, according to the shrinkage of the O⋯O distance, underlying and sizing the associated Ubbelhode effect.^{a} The conformation and structure of the complex, and an estimation of the Ubbelhode effect have been obtained by combining the experimental data with the results of MP2/6-311++G** ab initio calculations.

aA. R. Ubbelhode, K. J. Gallagher, Acta Crystallogr. , 1955, 8, 71